# Rephasing invariant formulae for general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles

### v1

@@ IV Summary
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}\right]$ for CP phases $\delta^{(\alpha i)}$ associated with the nine parameterizations $V^{(\alpha i)}$ of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column with trivial phases in $V^{(\alpha i)}$.
Furthermore, we find exact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ between the nine phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of the unitarity triangles. These twelve relations revealed intrinsic underlying orders connecting the CP phases and angles, which thought to be random and independent. We show that a relation from the previous work $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is immediately reproduced as a special case ($\alpha=1,i=2$) of the first sum rules. These results provide a general and concise framework between differences of angles and phases, and offer a valuable guide for future experimental tests of CP violation.

### v2

@@ IV Summary
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}\right]$ for CP phases $\delta^{(\alpha i)}$ associated with the nine parameterizations $V^{(\alpha i)}$ of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column with trivial phases in $V^{(\alpha i)}$. In practice, since cancellations occur between the numerator and the denominator, these are a necessary and sufficient set of irreducible fifth-order invariants constructed using the determinant.
Furthermore, we find exact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}$ and $\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ between the nine phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of the unitarity triangles. These relations revealed intrinsic underlying orders connecting the CP phases and angles, which seem to be random and independent. We show that a relation from the previous work $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is immediately reproduced as a special case ($\alpha=1,i=2$) of the first sum rules. These results provide a general and concise framework for differences between angles and phases, and offer a valuable guide for future experimental tests of CP violation.